OpenStudy (anonymous):

y varies directly as the square of x. When x = 2, y = 20. Find y when x = 4.

OpenStudy (anonymous):

Time to translate ... y = kx^2 is for the first sentence. Ready to substitute?

OpenStudy (anonymous):

substitiute what ?

OpenStudy (anonymous):

\[20 = k2^2 = 4k \rightarrow k = 20/4 = 5\]\[y = 5\times x^{2}\] x = 4... so y = 5 * 4^2 = 80

OpenStudy (anonymous):

With y = kx^2 substitute x = 2, y = 20